# Correlated helimagnetic configuration in a nonsymmorphic magnetic nodal semimetal

### v2

@@ Abstract
The Harper equations have $q$ eigenvalues for a given $(k_{x},k_{y},k_{z}^{0})$. The original band is split into $q$ subbands due to the Kondo coupling to the CNHMC, and each subband has a reduced magnetic Brillouin zone: $-\pi\leq k_{x},k_{y}\leq\pi$ and $-\pi/q\leq k_{z}^{0}\leq\pi/q$.
@@
Figure 2: (color online) The band structures of Weyl fermions with the CNHMC. $Q=2\pi/3$, $t=t^{\prime}=0.1eV$, and $k_{x}=k_{y}=0$. (a) $K_{0}=10meV$, $v_{0}=5meV$. (b) $K_{0}=200meV$, $v_{0}=5meV$. (c) $K_{0}=200meV$, $v_{0}=5meV$ with open boundary conditions of 50 sites along $y$-axis. (d) $K_{0}=10meV$, $v_{0}=50meV$. The red points in (a), (b), and (d) label the Weyl nodes. The red line in (c) stands for the Fermi arc.
Magnetic phase diagram. We turn to study the magnetic phase diagrams of the model Hamiltonian (1). For a fixed magnetic configuration, which is either the FM or the CNHMC, we compare the ground state energy per site at zero temperature and the free energy $f$ per site at finite temperature of Eq. (1). For the CNHMC, one takes $Q=2\pi/3$ and the FM is given by $Q=0$. To obtain the spectra, we use exact diagonalization in the calculations and choose periodic boundary conditions for a cubic lattice with $L_{x}=L_{y}=L_{z}=20$.
At zero temperature, we find that for a fixed $J_{0}$, the CNHMC is favored for large DM interaction $D_{0}$. The phase boundary moves downward when $K_{0}$ increases (see Fig. 3a). At finite temperature, by choosing $J_{0}=-3meV$, $D_{0}=-4meV$, $K_{0}=10meV$, with the other parameters the same as those in Fig. 2a, a phase transition occurs at about $T_{c}=4.5K$ (Fig. 3b), which is qualitatively consistent with experiments in Yu2023 , although the ordered phase is a canted magnetic order instead of the FM. By increasing $K_{0}$, $T_{c}$ also increases (see Fig. 3b).
Figure 3: (color online) The parameters are the same as in Fig. 2a except for $K_{0}$. $K_{0}=10meV$ for the red lines and $K_{0}=50meV$ for the blue lines. (a) The zero-temperature phase diagram with $J_{0}$ versus $D_{0}$. The gray dashed line stands for the case with $D_{0}=0$, for which the ground state energy per site is the same for both phases along the lines. Above (below) the lines, the FM (helical) configuration is energetically favored. (b) The free energy per site $f$ versus temperature $T$ for $J_{0}=-3meV$ and $D_{0}=-4meV$. The solid (dashed) lines stand for $f$ with the FM (helical) configuration.
CME with CNHMC. To further explore the effects of the CNHMC in this type of magnetic Weyl semimetals, we study the CME CME . For an isolated Weyl node given by the low-energy Hamiltonian $h_{w}=b_{0}+{\boldsymbol{\sigma}}\cdot({\bf k-b})$, where ${\bf b}$ and $b_{0}$ denote the shift in momentum and energy, respectively. Its response to the external magnetic field is described by the topological $\theta$-term, $\theta_{{\bf r},t}{\bf E\cdot B}$, with $\theta_{{\bf r},t}=2({\bf b\cdot r}-b_{0}t)$ Burkov2012 ; Yamamoto2012 ; Tewari2013 ; Grushin2012 , which originates from the chiral anomaly of the Weyl fermion Adler1969 ; Jackiw1969 ; no-go and corresponds to the charge density $\rho\propto{\bf b\cdot B}$ and current ${\bf j}\propto({\bf b\times E}-b_{0}{\bf B})$. The ${\bf b}$-related terms contribute to the anomalous Hall effect, while the $b_{0}$ term contributes to the CME which is related to the giant negative magnetoresistance Franz2013 ; Burkov2015 ; Spivak2013 ; chen2015 . In our model Hamiltonian (Correlated helimagnetic configuration in a nonsymmorphic magnetic nodal semimetal), $b_{0}\propto v_{0}$ and ${\bf b}\propto(0,K_{0},0)$ in the FM limit $Q=0$. For the CNHMC with $|Q|=2|k_{0}|$, ${\bf b}$ is not well-defined for small $K_{0}$ because the Weyl nodes are close to each other due to band-folding.
To explore the CME with the CNHMC, we apply a uniform magnetic field in the $z$-direction in the Hamiltonian (Correlated helimagnetic configuration in a nonsymmorphic magnetic nodal semimetal). In the $x$-$y$ plane, the lattice sites are labeled by $(n,m)$, and the gauge $\theta^{x}_{(n,m),(n+1,m)}=0$ and $\theta^{y}_{(n,m),(n,m+1)}=2\pi\phi n$ is used Kohmoto1989 ; Kohmoto1990 ; Hatsugai1993 . Under this gauge, the flux through a unit plaquette is $2\pi\phi$. The current arising from the CME reads Franz2013
where $n_{F}$ is the Fermi distribution function. The periodic boundary conditions in the $y$- and $z$-directions are used. Both periodic and open boundary conditions are considered in the $x$-direction with a length $L_{x}$. We set $\phi=1/5$.
Figure 4: (color online) The current $J_{z}$ of the CME at zero temperature with $t=t^{\prime}=0.1eV$, $\phi=1/5$, and $L_{x}=20$ is shown with periodic boundary conditions for the red solid lines and open boundary conditions for the blue dashed lines. $\Lambda=\pi/6$. (a) $K_{0}=10meV$, $Q=2\pi/3$; (b) $K_{0}=10meV$, $Q=0$; (c) $v_{0}=5meV$, $Q=2\pi/3$; (d) $v_{0}=5meV$, $Q=0$.
To see the CME in Weyl semimetals, a cutoff $\Lambda\ll\pi$ in the summation of $k_{z}$ is taken instead of summing over the whole Brillouin zone and the current grows linearly in $b_{0}$ for a Weyl semimetal Franz2013 . With this scenario, we calculate the CME with $L_{x}=20$, $\Lambda=\pi/6$, and plot $J_{z}$ by varying $v_{0}$ and $K_{0}$ at zero temperature in Fig. 4. We find that the Fermi arc states have a negative contribution to $J_{z}$ because $J_{z}$ under open boundary conditions is slightly smaller than that under periodic conditions. In the FM case, $J_{z}$ is linearly dependent on $v_{0}$ as expected (see Fig. 4b), while for the CNHMC, $J_{z}$ remains a constant when $v_{0}<3K_{0}$ and then grows (see Fig. 4a). For a large $v_{0}$, the folded Weyl nodes are separated far enough and isolated (see Fig. 2d). Therefore, the linear behavior of $J_{z}$ versus $v_{0}$ is expected. In the $J_{z}$ constant regime, $J_{z}$ decreases when $K_{0}$ increases. This shows that the Kondo coupling suppresses the CME with the CNHMC (see Fig. 4c) while it raises the CME with the FM order (see Fig. 4d). The key observation is that the CME in the CNHMC is about 20 times smaller than that in the FM order, which may explain the nearly zero behavior observed near $T=12.8K$ in CeAlGe Yu2023 .
Conclusions. We constructed a minimal lattice model with Kondo, FM, and DM interactions for the Weyl semimetal in which both $\Theta$ and $P$ symmetries are broken, which caught the key correlated physics in ReAlX. We found that the CNHMC induces a magnetic Brillouin zone in the band structure of the conduction electrons. The magnetic phase transitions from the ordered phase to the CNHMC were explored. We also calculated the CME with the CNHMC, which is about 20 times smaller than that with the FM order. This may provide an explanation for the anomalous transport phenomena and nearly zero magnetoresistance observed in CeAlGe.
The authors thank Long Liang and Yong-Shi Wu for useful discussions. This work is supported by the National Natural Science Foundation of China with Grant No. 12174067 (XL, YMZ, and YY).

### v3

@@ III.1 Band structures of the fermion
The Harper equations have $q$ eigenvalues for a given $(k_{x},k_{y},k_{z}^{0})$. The original band is split into $q$ subbands due to the Kondo coupling to the CNHMC, and each subband has a reduced magnetic Brillouin zone: $\pi\leq k_{x},k_{y}\leq\pi$ and $\pi/q\leq k_{z}^{0}\leq\pi/q$.
In ReAlX, the Fermi velocity is $\sim 1eV\mathring${A}, and the lattice constant is $\sim 10\mathring${A}, and then $t\sim 100meV$. The Curie temperature is about $30K$, and then $|J_{0}|\sim 2.5meV$. From the first-principles calculations, $|K_{0}|\sim|D_{0}|\sim|J_{0}|$ are of the same order Tafti2023 ; Tafti2021 ; Tafti2020 ; Broholm2021 ; Yamada2024 ; Li2023 ; Alam2023 ; Yu2023 ; Yang2023 ; Terashima2024 . With these parameters, we plot the band structures of $H_{0}$ (II.1) with the CNHMC in Fig. 2. The observed helimagnetic ordering in SmAlSi with nesting vector ${\bf Q}=\frac{2\pi}{3}(0.979,0.979,0)$ Tafti2023 motivates our choice of the commensurate $Q=2\pi/3$ ($1/3$ Brillouin zone length) as the Weyl node nesting vector. We find that, for the commensurate nesting $Q=2\pi/3$, the Weyl nodes are folded to the boundaries of the magnetic Brillouin zone for a small $K_{0}$ (see Fig. 2a). The monopole charges of the left and right Weyl nodes (red points) in Fig. 2b are 1-1 and +1+1, whereas the monopole charges of the other band crossings induced by the band folding of CNHMC are zero. This topological feature is also confirmed by Fermi arc states under open-boundary conditions in Fig. 2c. As $K_{0}$ increases, the Weyl nodes move towards each other (see Fig. 2b), which may explain the slightly incommensurate behavior of the helimagnetic configuration observed in experiments Tafti2023 . For a strong enough $K_{0}$, the Weyl nodes merge together and finally open a gap. The Fermi arcs that connect the Weyl nodes will also evolve from shortening to disappearing; see Fig. 2c (for more details, see Appendix C). These phenomena reflect the competition between the monopole topology of the Weyl node in momentum space and the trivial topology of the helimagnet in real space. The $P$-breaking term $v_{0}$ shifts the energy of every Weyl node and does not alter the location of the Weyl nodes (see Fig. 2d).
Figure 2: (color online) The band structures of Weyl fermions with CNHMC for parameters $Q=2\pi/3$, $t=t^{\prime}=0.1$ eV, and $k_{x}=k_{y}=0$. (a) $K_{0}=10$ meV. (b) $K_{0}=200$ meV. (c) Open boundary conditions for (b) with 50 sites along the $y$ axis. (d) $K_{0}=10$ meV. Parameter values: (a-c) $v_{0}=5$ meV and (d) $v_{0}=50$ meV. The red points in (a), (b), and (d) indicate the Weyl nodes. The red line in (c) represents the Fermi arc.
III.2 Magnetic phase diagram
We turn to study the magnetic phase diagrams of the model Hamiltonian (1). For a fixed magnetic configuration, which is either the ferromagnet or the CNHMC, we compare the ground state energy per site at zero temperature and the free energy $f$ per site at finite temperature of Eq. (1). For the CNHMC, one takes $Q=2\pi/3$ and the ferromagnet is given by $Q=0$. To obtain the spectra, we use exact diagonalization in the calculations and choose periodic boundary conditions for a cubic lattice with $L_{x}=L_{y}=L_{z}=20$.
At zero temperature, we find that for a fixed $J_{0}$, the CNHMC is favored for large DM interaction $D_{0}$. The phase boundary moves downward when $K_{0}$ increases (see Fig. 3a). At finite temperature, by choosing $J_{0}=-3meV$, $D_{0}=-4meV$, $K_{0}=10meV$, with the other parameters the same as those in Fig. 2a, a phase transition occurs at about $T_{c}=4.5K$ (Fig. 3b), which is qualitatively consistent with experiments in Ref. Yu2023 , although the ordered phase is a canted magnetic order instead of the ferromagnet. By increasing $K_{0}$, $T_{c}$ also increases (see Fig. 3b).
Figure 3: (color online) The parameters are the same as in Fig. 2a except for $K_{0}$. $K_{0}=10meV$ for the red lines and $K_{0}=50meV$ for the blue lines. (a) The zero-temperature phase diagram with $J_{0}$ vs $D_{0}$. The gray dashed line stands for the case with $D_{0}=0$, for which the ground state energy per site is the same for both phases along the lines. Above (below) the lines, the FM (helical) configuration is energetically favored. (b) The free energy per site $f$ vs temperature $T$ for $J_{0}=-3meV$ and $D_{0}=-4meV$. The solid (dashed) lines stand for $f$ with the FM (helical) configuration.
IV CME with CNHMC
To further explore the effects of the CNHMC in this type of magnetic Weyl semimetals, we study the CME CME . For an isolated Weyl node given by the low-energy Hamiltonian $h_{w}=b_{0}+{\boldsymbol{\sigma}}\cdot({\bf k-b})$, where $\bf b$} and $b_{0}$ denote the shift in momentum and energy, respectively. Its response to the external magnetic field is described by the topological $\theta$ term, $\theta_{{\bf r},t}{\bf E\cdot B}$, with $\theta_{{\bf r},t}=2({\bf b\cdot r}-b_{0}t)$ Burkov2012 ; Yamamoto2012 ; Tewari2013 ; Grushin2012 , which originates from the chiral anomaly of the Weyl fermion Adler1969 ; Jackiw1969 ; no-go and corresponds to the charge density $\rho\propto{\bf b\cdot B}$ and current ${\bf j}\propto({\bf b\times E}-b_{0}{\bf B})$. The $\bf b$}-related terms contribute to the anomalous Hall effect, while the $b_{0}$ term contributes to the CME which is related to the giant negative magnetoresistance Franz2013 ; Burkov2015 ; Spivak2013 ; chen2015 . In our model Hamiltonian (II.1), $b_{0}\propto v_{0}$ and ${\bf b}\propto(0,K_{0},0)$ in the FM limit $Q=0$. For the CNHMC with $|Q|=2|k_{0}|$, $\bf b$} is not well defined for small $K_{0}$ because the Weyl nodes are close to each other due to band folding.
To explore the CME with the CNHMC, we apply a uniform magnetic field in the $z$ direction in the Hamiltonian (II.1). In the $x$-$y$ plane, the lattice sites are labeled by $(n,m)$, and the gauge $\theta^{x}_{(n,m),(n+1,m)}=0$ and $\theta^{y}_{(n,m),(n,m+1)}=2\pi\phi n$ is used Kohmoto1989 ; Kohmoto1990 ; Hatsugai1993 . Under this gauge, the flux through a unit plaquette is $2\pi\phi$. The current arising from the CME reads Franz2013
where $n_{F}$ is the Fermi distribution function. The periodic boundary conditions in the $y$ and $z$ directions are used. Both periodic and open boundary conditions are considered in the $x$ direction with a length $L_{x}$. We set $\phi=1/5$.
Figure 4: (color online) The current $J_{z}$ of the CME at zero temperature with $t=t^{\prime}=0.1eV$, $\phi=1/5$, and $L_{x}=20$ is shown with periodic boundary conditions for the red solid lines and open boundary conditions for the blue dashed lines. $\Lambda=\pi/6$. (a) $K_{0}=10meV$, $Q=2\pi/3$; (b) $v_{0}=5meV$, $Q=2\pi/3$; (c) $K_{0}=10meV$, $Q=0$; (d) $v_{0}=5meV$, $Q=0$.
To see the CME in Weyl semimetals, a cutoff $\Lambda\ll\pi$ in the summation of $k_{z}$ is taken instead of summing over the whole Brillouin zone and the current grows linearly in $b_{0}$ for a Weyl semimetal Franz2013 . With this scenario, we calculate the CME with $L_{x}=20$, $\Lambda=\pi/6$, and plot $J_{z}$ by varying $v_{0}$ and $K_{0}$ at zero temperature in Fig. 4. We find that the Fermi arc states have a negative contribution to $J_{z}$ because $J_{z}$ under open boundary conditions is slightly smaller than that under periodic conditions. In the FM case, $J_{z}$ is linearly dependent on $v_{0}$ as expected (see Fig. 4c), while for the CNHMC, $J_{z}$ remains a constant when $v_{0}<3K_{0}$ and then grows (see Fig. 4a). For a large $v_{0}$, the folded Weyl nodes are separated far enough and isolated (see Fig. 2d). Therefore, the linear behavior of $J_{z}$ versus $v_{0}$ is expected. In the $J_{z}$ constant regime, $J_{z}$ decreases when $K_{0}$ increases. This shows that the Kondo coupling suppresses the CME with the CNHMC (see Fig. 4b) while it raises the CME with the FM order (see Fig. 4d). The key observation is that the CME in the CNHMC is about 20 times smaller than that in the FM order, which may explain the nearly zero behavior observed near $T=12.8K$ in CeAlGe Yu2023 .
V Conclusions
We constructed a minimal lattice model with Kondo, FM, and DM interactions for the Weyl semimetal in which both $\Theta$ and $P$ symmetries are broken, which caught the key correlated physics in ReAlX. We found that the CNHMC induces a magnetic Brillouin zone in the band structure of the conduction electrons. The magnetic phase transitions from the ordered phase to the CNHMC were explored. We also calculated the CME with the CNHMC, which is about 20 times smaller than that with the FM order. This may provide an explanation for the anomalous transport phenomena and nearly zero magnetoresistance observed in CeAlGe.
Acknowledgements.
The authors thank Long Liang and Yong-Shi Wu for useful discussions. This work is supported by the National Natural Science Foundation of China with Grant No. 12174067 (XL, YMZ, and $Y$).
DATA AVAILABILITY
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
Appendix A Continuum limit and exactly soluble correlated helimagnet
When $|Q|\ll 2|k_{0}|$, the low energy effective Hamiltonian for the two Weyl nodes becomes
$H_{eff}=\int d^{3}r\sum_{\nu=\pm}(-iv_{\nu}\psi^{\dagger}_{\nu}{\boldsymbol{\sigma}}\cdot\nabla\psi_{\nu}+{K_{0}}{\bf M}\cdot{\bf s_{\nu}}+{\nu b_{0}})+$..., (8)
where $\nu$ labels the valley of the Weyl nodes, $v_{\nu}$ is the Fermi velocity of the Weyl node $\nu$, and $b_{0}$ is the energy offset between the two Weyl nodes caused by the inversion symmetry breaking. The ”…” terms include the higher-order terms of derivatives and the exchange terms between $\bf M$}s. Due to the energy offset $b_{0}$ between the two Weyl nodes in a non-centrosymmetric system, we can integrate out one Weyl fermion that is below the chemical potential, say, $\psi_{-}$. The effective Hamiltonian is given by (the constant $b_{0}$ is omitted) Yu16 ; Yu2023
$H_{eff}$ $=$ $\int d^{3}r\biggl{[}-i(\psi^{\dagger}\boldsymbol{\sigma}\cdot\nabla\psi+K_{0}{\bf M}\cdot{\bf s})$ (9)
$+\frac{D}{4}{\bf M}\cdot{\nabla\times\bf M}+\frac{J}{2}(\nabla{\bf M})^{2}\biggr{]},$
where the suffix + is omitted; $J\propto K_{0}^{2}-2J_{0}$ and $D\propto v_{-}K_{0}^{2}+4D_{0}$ Xiao2015 ; rkky2 . We have set $\hbar=v_{+}=1$ and neglected the high-order terms. When $K_{0}=0$, the ground state configuration of the normalized $\bf M$} is an uncorrelated helimagnet, ${\bf M}=[{\bf e}_{1}\sin({\bf Q}\cdot{\bf r}+\varphi_{0})\pm{\bf e}_{2}\cos({\bf Q}\cdot{\bf r}+\varphi_{0})]$ with $Q=|{\bf Q}|=|D/2J|$ and ${\bf Q}\propto{\bf e}_{1}\times{\bf e}_{2}$ Nagaosa2012 . When $K_{0}\neq 0$, the spin $\bf s$} and $\bf M$} are coupled. The equation of motion for $\psi$ and the ground state configuration of $\bf M$} satisfy Yu16
$-i\boldsymbol{\sigma}\cdot(\nabla+i\frac{K_{0}}{2}{\bf M})\psi=E_{e}\psi,$ (10)
$K_{0}{\bf s}+\frac{D}{2}\nabla\times{\bf M}-{J}\nabla^{2}{\bf M}=0.$ (11)
Consider the following ansatz Don ; freund ; adam ; sol1 ; sol2 ; sol3 ; 3T :
${\bf M}$ $=$ $(\sin[Qx],\cos[Qx],0),$ (12)
${\bf B}_{0}$ $=$ $(\pm)\frac{K_{0}}{2}\nabla\times{\bf M},$ (13)
@@ Appendix A Continuum limit and exactly soluble correlated helimagnet
One can check ${\bf B}_{0}=\chi^{\dagger}{\boldsymbol{\sigma}}\chi$ and $B_{0}=\chi^{\dagger}\chi$. Therefore this provides a map from $SU(2)$ to $SO(3)$. Then the equation of motion for $\psi=\chi/\sqrt{\chi^{\dagger}\chi}$ reduces to
$(\frac{Q}{2}+\frac{K_{0}}{2})\psi=E_{e}\psi,$ (17)
and the ground state equation for $\bf M$} reduces to
$K_{0}{\bf M}+\frac{DQ}{2}{\bf M}-|J|Q^{2}{\bf M}=0.$ (18)
Therefore one obtains
$Q=\frac{D\pm\sqrt{D^{2}+16|J|K_{0}}}{4|J|}.$ (19)
We shall emphasize that although the helical configuration here has the same form as the case for $K_{0}=0$, $|{\bf Q}|$ is correlated with the Weyl fermions via the Kondo coupling. It can be proved that the ground state solution can be found for a general helical magnetic configuration ${\bf M}=\frac{|Q|}{K_{0}}({\bf e}_{1}\sin({\bf Q\cdot r}+\phi_{0})\pm{\bf e}_{2}\cos({\bf Q\cdot r}+\phi_{0}))$ Yu16 .
We present a more concrete solution with the ansatz ${\bf M}=\frac{Q}{K}(\sin[Qz],\cos[Qz],0)$. Assuming the periodic boundary condition, the spectrum $E$ of the fermions is determined by solving the equation of motion under the effective Hamiltonian (10):
$(-i\boldsymbol{\sigma}\cdot\nabla+\boldsymbol{\sigma}\cdot\frac{K}{2}{\bf M})\psi=E\psi.$ (20)
In order to find the zero-energy solution, we focus along one direction, say, the $z$ axis, and take the following ansatz for the wave function:
$\psi_{0}({\bf r})=\psi_{n}(z)$
$\propto e^{i(k_{z}+nQ)z}\left(\begin{array}[]{c}1\\ -\frac{2i}{K}(k_{z}+nQ-E)e^{-iQz}\end{array}\right),$ (23)
where $k_{z}$ is the wave vector along the $z$ axis, $Q$ is the magnetic wave vector of $\bf M$} , and $n$ labels the $n$th magnetic Brillouin zone determined by $Q$. Then, through a direct calculation, the spectrum reads
$E_{\pm}=\frac{Q}{2}\pm\sqrt{(k_{z}+(n+\frac{1}{2})Q)^{2}+\frac{K^{2}}{4}}.$ (24)
The zero energy occurs when $|Q|=|K|$ (which is consistent with the normalization $|{\bf M}|=1$), and $k_{z}=-(n+\frac{1}{2})Q$. The diagram of the spectra (24) is shown in Fig. 1 of the main text. The spin of these zero modes is the same, i.e., ${\bf s}_{n}$${\bf s}=\frac{1}{2}\frac{Q}{K}{\bf M}$. Eq. (11) determines $Q$ as
$\frac{Q}{2K}+\frac{DQ}{2K}-\frac{|J|Q^{2}}{K}=0\Rightarrow Q=\frac{1}{2|J|}(1+D),$ (25)
which also gives a constraint between $K$ and $J/D$. For a pinned helimagnet with $Q>0$, say, in the $z$ direction, the dispersion near $k_{z}=-(n+\frac{1}{2})Q$ for a given zero mode reads,
$E_{-}(k_{z})\approx-\frac{1}{|K|}(k_{z}+(n+\frac{1}{2})Q)^{2}\equiv-\frac{1}{|K|}q_{z}^{2}.$ (26)
We now perform the perturbation by taking $H_{1}=k_{x}\sigma^{x}+k_{y}\sigma^{y}$ because for the pinned helimagnet, the Hamiltonian (9) in the main text is translationally invariant in the $x$-$y$ plane. The first-order perturbation is zero. The matrix elements between
$\psi_{n-}(z)\propto e^{i(k_{z}+nQ)z}\left(\begin{array}[]{c}1\\ -\frac{i2}{K}(k_{z}+nQ-E_{n-})e^{-iQz}\end{array}\right)$ (29)
$=$ $e^{i(k_{z}+nQ)z}\left(\begin{array}[]{c}1\\ A_{n}e^{-iQz}\end{array}\right)$ (32)
and $\psi_{n^{\prime}-}(z)$ are
$\langle\psi_{n^{\prime}-}(z)|H_{1}|\psi_{n-}(z)\rangle$ (33)
$=$ $\delta_{n,n^{\prime}+1}A_{n}(k_{x}-ik_{y})+\delta_{n,n^{\prime}-1}A_{n^{\prime}}(k_{x}+ik_{y}).$
The energy differences at $k_{z}=-(n+\frac{1}{2})Q$ are
$E_{-}(n,k_{z})-E_{-}(n\pm 1,k_{z})=\frac{\sqrt{5}-1}{2}K,$ (34)
The second order perturbation gives
$E^{(2)}_{-}$ $=$ $\sum_{n^{\prime}\neq n}\frac{\langle\psi_{n-}(z)|H_{1}|\psi_{n^{\prime}-}(z)\rangle\langle\psi_{n^{\prime}-}(z)|H_{1}|\psi_{n-}(z)\rangle}{E_{n}-E_{n^{\prime}}}$ (35)
$=$ $\frac{2}{\sqrt{5}-1}\frac{k_{x}^{2}+k_{y}^{2}}{K}.$
Thus, the dispersions near $k_{z}=-(n+1/2)Q$ are
$E_{-}({\bf k})$ $=$ $-\frac{1}{K}q_{z}^{2}+\frac{2}{(\sqrt{5}-1)K}k_{t}^{2}$ (36)
$\equiv$ $-\frac{q_{z}^{2}}{2m_{z}}+\frac{k_{t}^{2}}{2m_{t}}=-\varepsilon_{z}+\varepsilon_{t},$
with ${\bf k}_{t}=(k_{x},k_{y})$. The dispersion is of a saddle shape (see Fig. 5); namely, in the $z$ direction, it is hole-type, while in the $x$-$y$ plane, it is particle-type. This means that the dynamics parallel to the helimagnet and perpendicular to the helimagnet are separated.
Figure 5: (color online) A schematic diagram for the dispersion of $E_{-}$ (36) near the blue points in Fig. 1. $K=1$. The dispersion has a saddle shape. It is a parabola along $k_{t}$ and a downward one along $q_{z}$.
Appendix B A brief review of the two-dimensional lattice model in a constant external magnetic field and the Harper equation
Consider a tight-binding Hamiltonian on the square lattice in a magnetic field Kohmoto1989 ; Kohmoto1990 ; Hatsugai1993 :
$H$ $=$ $-t_{a}\sum_{\langle i,j\rangle}c_{x,j}^{\dagger}c_{x,i}e^{i\theta^{x}_{ij}}$ (37)
$-t_{b}\sum_{\langle i,j\rangle}c_{y,j}^{\dagger}c_{y,i}e^{i\theta^{y}_{ij}}+h.c.,$
where $\theta_{ij}$ is the phase factor of the gauge field which satisfies $\sum_{\boxdot}\theta_{ij}=2\pi\phi$ in a unit cell, and $\phi=p/q$ with $p,q$ coprime. $t_{a}$ and $t_{b}$ are the nearest-neighbor hopping constants. A site $i$ on the square lattice has Cartesian coordinates $(n,m)$ where $n$ and $m$ are integers. We choose the gauge $\theta^{x}_{ij}=0$, and $\theta^{y}_{ij}=2\pi\phi n$ for the link between $i=(n,m)$ and $j=(n,m+1)$. This gauge gives a uniform magnetic field the flux of which through a plaquette is $2\pi\phi$. In the momentum space,
$H({\bf k})$ $=$ $-2t_{a}\cos k_{x}c^{\dagger}(k_{x},k_{y})c(k_{x},k_{y})$ (38)
$-t_{b}[e^{-ik_{y}}c^{\dagger}(k_{x}+2\pi\phi,k_{y})c(k_{x},k_{y})$
$+e^{ik_{y}}c^{\dagger}(k_{x}-2\pi\phi,k_{y})c(k_{x},k_{y})],$
where $\pi\leq k_{x},k_{y}\leq\pi$. Since $k_{x}$ is coupled to $k_{x}\pm 2\pi\phi$, the Schrödinger equation $H|\Psi\rangle=E|\psi\rangle$ is reduced to the Harper equation:
$-t_{b}(e^{-ik_{y}}\psi_{j-1}+e^{ik_{y}}\psi_{j+1})-2t_{a}\cos(k^{0}_{x}+2\pi\phi j)\psi_{j}$
$=E(k_{x}^{0},k_{y})\psi_{j},$ (39)
where $k_{x}=k_{x}^{0}+2\pi\phi j$ and $\phi_{j+q}=\phi_{j}$. And the state $|\Psi\rangle$ is
$|\Psi\rangle=\sum_{j=1}^{q}\psi_{j}c^{\dagger}(k_{x}^{0}+2\pi\phi j,k_{y})|0\rangle.$ (40)
Appendix C Evolution of Weyl nodes with respect to the Kondo coupling $K_{0}$
We plot the evolution of Weyl nodes with respect to the Kondo coupling $K_{0}$ in Fig. 6. We find that by increasing the Kondo coupling, the Weyl nodes move closer to each other. Finally, they annihilate and open a gap. This process reflects the competition between the monopole topology of the Weyl node and the trivial topology of the helimagnet via Kondo coupling $K_{0}$. For a small $K_{0}$, the Weyl node structure remains stable, while for a large $K_{0}$, the conducting fermions acquire a gap and become trivial.
@@ Appendix C Evolution of Weyl nodes with respect to the Kondo coupling $K_{0}$
Figure 6: (Color online) The evolution of Weyl nodes. The parameters are chosen as $Q=2\pi/3$, $t=t^{\prime}=0.1eV$, and $k_{x}=k_{y}=0$. (a) $K_{0}=250meV$, (c) $K_{0}=290meV$, (e) $K_{0}=350meV$. (b), (d), and (f) are the corresponding band structures with open boundary conditions for 50 sites along the $y$-direction. The red points are Weyl nodes, and the red lines are the Fermi arcs.